# GenPO: Generative Diffusion Models Meet On-Policy Reinforcement Learning

### v2

@@ 4 GenPO: An On-policy RL Method based on Diffusion Models
As mentioned in Section 3.2, the log-likelihood of the given samples cannot be computed directly in generative models. However, to integrate diffusion policy into on-policy RL, we must present the explicit formula of log-likelihood under the given state-action pair for the policy update. Hence, we first elucidate the intrinsic reasons that make the computation of the log-likelihood in diffusion models challenging. Then, we address these challenges by (1) applying the exact diffusion inversion technique wallace2023edict and utilizing the change of variables technique like normalizing flow rezende2015variational , (2) constructing a reformulated Markov decision process with a modified action space, along with a trick that maps the generated actions back to the original action space.
Besides, based on this framework, we also introduce an unbiased estimation method for both the entropy of the diffusion policy and the Kullback-Leibler (KL) divergence between different generative diffusion policies. Notably, this has not yet been accomplished by previous diffusion-based RL methods. In addition, since policy update is actually performed on a reformulated MDP, extraneous exploration is induced, thereby degrading RL learning efficiency. To address this issue, we also introduce an auxiliary compression loss term. Leveraging these techniques, we finally present the practical implementation of the proposed GenPO algorithm, which successfully incorporates the generative policy to PPO schulman2017proximal in an effective manner. Figure 2 shows the training and inference process of the proposed GenPO.
@@ 4.1 Challenges in Calculating Diffusion Probability
Recalling existing generative models, such as VAE kingma2013auto , GAN creswell2018generative , normalizing flow rezende2015variational , and diffusion model song2020score , it can be found that only normalizing flow and its variants like flow GAN grover2017flow allow the exact likelihood computation via the change of variables theorem rezende2015variational . In contrast, for other generative models, probability densities or related statistics can only be obtained approximately via special designs ding2024diffusion ; celik2025dime ; wang2024diffusion . Applying such approximate probability density estimates for policy update is unacceptable in on-policy RL algorithms.
@@
Lemma 1. (Change of Variables  blitzstein2019introduction ) Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is an invertible and smooth mapping. If we have the random variable $X\sim q(x)$ and the random variable $Y=f(X)$ transformed by function $f$, the distribution $p(y)$ of $Y$ is
@@
However, as mentioned in wallace2023edict ; wang2024belm ; zhang2024exact , it is nontrivial to realize an exact invertible diffusion model. This is because of the inconsistency between the forward and reverse processes of the generative diffusion model. (6) shows the forward and reverse process of DDIM song2020denoising . It can be observed that this inconsistency is caused by the approximation of $x_{t}$ with $x_{t-1}$ in the forward process, since the $x_{t}$ is unavailable in the forward step $t-1$.
@@ 4.2 Exact Diffusion Inversion in Reformulated MDP
Consequently, to realize an invertible diffusion model, this issue must be addressed. Motivated by EDICT wallace2023edict , we realize the invertible diffusion model via maintaining two coupled noise vectors and updating them alternately in the forward and reverse processes of the generative diffusion model, and then extend it into flow matching lipmanflow ; liu2022flow for fast generation as shown in (7, 8). However, the two coupled noise vectors also lead to a doubled sample space. This implies we cannot directly apply this technique to diffusion policy when the dimension of the action space is odd.
@@
$\begin{array}[]{lll}\text{Reverse:}&\tilde{x}_{t+\Delta t}=x_{t}+v_{\theta}(y_{t},t)\Delta t,&\tilde{y}_{t+\Delta t}=y_{t}+v_{\theta}(\tilde{x}_{t+\Delta t},t)\Delta t\\ \text{Mixing:}&x_{t+\Delta t}=p\cdot\tilde{x}_{t+\Delta t}+(1-p)\cdot\tilde{y}_{t+\Delta t},&y_{t+\Delta t}=p\cdot\tilde{y}_{t+\Delta t}+(1-p)\cdot x_{t+\Delta t}\\ \end{array}$ (7)
@@
$\begin{array}[]{lll}\text{Unmixing:}&\tilde{y}_{t+\Delta t}=\frac{y_{t+\Delta t}-(1-p)x_{t+\Delta t}}{p}&\tilde{x}_{t+\Delta t}=\frac{x_{t+\Delta t}-(1-p)\tilde{y}_{t+\Delta t}}{p}\\ \text{Forward:}&y_{t}=\tilde{y}_{t+\Delta t}-v_{\theta}(\tilde{x}_{t+\Delta t},t)\Delta t,&x_{t}=\tilde{x}_{t+\Delta t}-v_{\theta}(y_{t},t)\Delta t\par\end{array}$ (8)
@@ 4.3 Practical Implementation
To enhance GenPO’s exploration and stabilize training, we estimate policy entropy and introduce an adaptive KL-divergence-based learning rate schedule for flow policies. Unlike Gaussian policies, flow policies lack closed-form expressions for entropy and KL divergence, preventing direct computation. As a result, prior diffusion RL algorithms ding2024diffusion ; wang2024diffusion ; celik2025dime have relied on heuristic approximations to encourage exploration. In contrast, GenPO is exactly invertible (Section 4.2), allowing precise computation of the log-likelihood of any state-action pair via the change-of-variables (5) in Section 4.1. This enables unbiased estimation of both entropy and KL divergence. Accordingly, we define the entropy loss in Eq. 9, which also incorporates importance sampling to account for the mismatch between the behavior and target policies.
$\mathcal{L}^{ENT}(\pi_{\theta}):=\mathbb{E}_{s,\tilde{a}\sim\pi_{\theta_{old}}}\left[\frac{\pi_{\theta}(\tilde{a}\mid s)}{\pi_{\theta_{old}}(\tilde{a}\mid s)}\log\left(\pi_{\theta}(\tilde{a}\mid s)\right)\right]$ (9)
Likewise, an unbiased estimation of the KL divergence can be obtained, thereby enabling dynamic adjustment of the learning rate. Finally, the practical formula for computing the KL divergence between diffusion policies is
@@
5:         Use $a=\frac{x+y}{2}$ to interact with the environment for $N$ timesteps
@@ 5.1 Comparative Evaluation
DDPG lillicrap2015continuous 62.96(5.18) 63.34(7.39) -1.61(0.76) 0.03(0.13) -0.86(0.55) 9.09(0.81) -6.63(1.04) 6209.59(559.44)
TD3 fujimoto2018addressing 67.80(7.05) 82.36(4.70) 0.03(0.12) 0.17(0.15) -0.89(0.77) 8.44(1.05) -4.97(0.87) 9386.64(314.37)
SAC haarnoja2018soft2 38.68(21.84) 49.52(16.89) -1.10(0.95) 38.46(22.20) -5.12(0.14) 6.67(2.02) -5.05(0.18) 8459.74(135.79)
DACER wang2024diffusion 0.29(0.98) 0.58(0.01) -3.43(2.82) 25.39(22.99) -1.62(0.09) 6.37(2.07) -6.21(0.17) -207.43(88.82)
QVPO ding2024diffusion 7.19(2.32) 10.59(6.30) -0.12(0.14) 63.99(45.42) -1.81(0.36) 7.34(0.62) -4.72(0.16) 134.36(39.05)
PPO schulman2017proximal 146.94(10.61) 197.25(18.26) 78.14(50.43) 99.08(13.49) 9.80(4.78) 15.67(1.92) 18.97(2.02) 8402.21(435.64)
@@ 5.2 Ablation Study
Figure 4: Ablation study results. (a) Effect of varying the compression loss coefficient $\nu$ on training stability and final performance. (b) Impact of entropy and KL divergence estimation on exploration and convergence. (c) Performance under different mixing coefficients $p$ in flow policies.
@@
Entropy and KL Divergence. Figure 4(b) highlights the impact of our entropy and KL divergence estimation components. KL divergence estimation enables adaptive adjustment of the update learning rate, leading to faster and more stable convergence. In parallel, entropy estimation enhances exploration, resulting in significantly higher returns compared to variants without it. Experiments have shown that our estimation methods (section 4.3) for entropy and divergence are beneficial to the results. We adaptively adjust the learning rate through KL divergence on all tasks and add an entropy regularization term to the loss(12).
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (62303319, 62406195), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
@@ B.3 Hyperparameters
Learning rate for actor $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$
Learning rate for critic $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$
@@
Memory size N/A N/A $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$
@@
Memory size N/A N/A $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$
@@
Learning rate for actor $1\times 10^{-4}$ $1\times 10^{-4}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Learning rate for critic $1\times 10^{-4}$ $1\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
@@
Memory size N/A N/A $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$
@@
Memory size N/A N/A $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$
@@
Learning rate for actor $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Learning rate for critic $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
@@
Memory size N/A N/A $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$
@@
Learning rate for actor $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Learning rate for critic $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
@@
Memory size N/A N/A $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$
@@
Learning rate for actor $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Learning rate for critic $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
@@
Memory size N/A N/A $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$
@@
Memory size N/A N/A $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$ $1\times 10^{6}$

### v3

@@ 4 GenPO: An On-policy RL Method based on Diffusion Models
As mentioned in Section 3.2, the log-likelihood of the given samples cannot be computed directly in generative models. However, to integrate diffusion policy into on-policy RL, we must present the explicit formula of log-likelihood under the given state-action pair for the policy update. Hence, we first elucidate the intrinsic reasons that make the computation of the log-likelihood in diffusion models challenging. Then, we address these challenges by (1) applying the exact diffusion inversion technique Wallace et al. (2023) and utilizing the change of variables technique like normalizing flow Rezende and Mohamed (2015), (2) constructing a reformulated Markov decision process with a modified action space, along with a trick that maps the generated actions back to the original action space.
Besides, based on this framework, we also introduce an unbiased estimation method for both the entropy of the diffusion policy and the Kullback-Leibler (KL) divergence between different generative diffusion policies. Notably, this has not yet been accomplished by previous diffusion-based RL methods. In addition, since policy update is actually performed on a reformulated MDP, extraneous exploration is induced, thereby degrading RL learning efficiency. To address this issue, we also introduce an auxiliary compression loss term. Leveraging these techniques, we finally present the practical implementation of the proposed GenPO algorithm, which successfully incorporates the generative policy to PPO Schulman et al. (2017) in an effective manner. Figure 2 shows the training and inference process of the proposed GenPO.
@@ 4.1 Challenges in Calculating Diffusion Probability
Recalling existing generative models, such as VAE Kingma and Welling (2013), GAN Creswell et al. (2018), normalizing flow Rezende and Mohamed (2015), and diffusion model Song et al. (2020b), it can be found that only normalizing flow and its variants like flow GAN Grover et al. (2017) allow the exact likelihood computation via the change of variables theorem Rezende and Mohamed (2015). In contrast, for other generative models, probability densities or related statistics can only be obtained approximately via special designs Ding et al. (2024); Celik et al. (2025); Wang et al. (2024b). Applying such approximate probability density estimates for policy update is unacceptable in on-policy RL algorithms.
@@
Lemma 1. (Change of Variables  Blitzstein and Hwang (2019)) Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is an invertible and smooth mapping. If we have the random variable $X\sim q(x)$ and the random variable $Y=f(X)$ transformed by function $f$, the distribution $p(y)$ of $Y$ is
@@
However, as mentioned in Wallace et al. (2023); Wang et al. (2024a); Zhang et al. (2024), it is nontrivial to realize an exact invertible diffusion model. This is because of the inconsistency between the forward and reverse processes of the generative diffusion model. (6) shows the forward and reverse process of DDIM Song et al. (2020a). It can be observed that this inconsistency is caused by the approximation of $x_{t}$ with $x_{t-1}$ in the forward process, since the $x_{t}$ is unavailable in the forward step t1t-1.
@@ 4.2 Exact Diffusion Inversion in Reformulated MDP
Consequently, to realize an invertible diffusion model, this issue must be addressed. Motivated by EDICT Wallace et al. (2023), we realize the invertible diffusion model via maintaining two coupled noise vectors and updating them alternately in the forward and reverse processes of the generative diffusion model, and then extend it into flow matching Lipman et al. (2023); Liu et al. (2022) for fast generation as shown in (7, 8). However, the two coupled noise vectors also lead to a doubled sample space. This implies we cannot directly apply this technique to diffusion policy when the dimension of the action space is odd.
@@
[…]
@@
$\begin{array}[]{lll}\text{Unmixing:}\quad&\tilde{y}_{t+\Delta t}=\frac{y_{t+\Delta t}-(1-p)x_{t+\Delta t}}{p}&\tilde{x}_{t+\Delta t}=\frac{x_{t+\Delta t}-(1-p)\tilde{y}_{t+\Delta t}}{p}\\ \text{Forward:}\quad&y_{t}=\tilde{y}_{t+\Delta t}-v_{\theta}(\tilde{x}_{t+\Delta t},t)\Delta t,&x_{t}=\tilde{x}_{t+\Delta t}-v_{\theta}(y_{t},t)\Delta t\par\end{array}$ (8)
@@ 4.3 Practical Implementation
To enhance GenPO’s exploration and stabilize training, we estimate policy entropy and introduce an adaptive KL-divergence-based learning rate schedule for flow policies. Unlike Gaussian policies, flow policies lack closed-form expressions for entropy and KL divergence, preventing direct computation. As a result, prior diffusion RL algorithms Ding et al. (2024); Wang et al. (2024b); Celik et al. (2025) have relied on heuristic approximations to encourage exploration. In contrast, GenPO is exactly invertible (Section 4.2), allowing precise computation of the log-likelihood of any state-action pair via the change-of-variables (5) in Section 4.1. This enables unbiased estimation of both entropy and KL divergence. Accordingly, we define the entropy loss in Eq. 9.
$\mathcal{L}^{ENT}(\pi_{\theta}):=\mathbb{E}_{s,\tilde{a}\sim\pi_{\theta}}\left[\log\left(\pi_{\theta}(\tilde{a}\mid s)\right)\right].$ (9)
Besides, to enable adaptive scheduling of the learning rate, we also present an unbiased estimation of the KL divergence as shown in Algorithm 1, and the estimation formula for computing the KL divergence between diffusion policies is
@@
5:   Use $a=\frac{x+y}{2}$ to interact with the environment for NN timesteps
@@ 5.1 Comparative Evaluation
DDPG Lillicrap et al. (2015) 62.96(5.18) 63.34(7.39) -1.61(0.76) 0.03(0.13) -0.86(0.55) 9.09(0.81) -6.63(1.04) 6209.59(559.44)
TD3 Fujimoto et al. (2018) 67.80(7.05) 82.36(4.70) 0.03(0.12) 0.17(0.15) -0.89(0.77) 8.44(1.05) -4.97(0.87) 9386.64(314.37)
SAC Haarnoja et al. (2018b) 38.68(21.84) 49.52(16.89) -1.10(0.95) 38.46(22.20) -5.12(0.14) 6.67(2.02) -5.05(0.18) 8459.74(135.79)
DACER Wang et al. (2024b) 0.29(0.98) 0.58(0.01) -3.43(2.82) 25.39(22.99) -1.62(0.09) 6.37(2.07) -6.21(0.17) -207.43(88.82)
QVPO Ding et al. (2024) 7.19(2.32) 10.59(6.30) -0.12(0.14) 63.99(45.42) -1.81(0.36) 7.34(0.62) -4.72(0.16) 134.36(39.05)
PPO Schulman et al. (2017) 146.94(10.61) 197.25(18.26) 78.14(50.43) 99.08(13.49) 9.80(4.78) 15.67(1.92) 18.97(2.02) 8402.21(435.64)
@@ 5.2 Ablation Study
Figure 4: Ablation study results. (a) Effect of varying the compression loss coefficient $\nu$ on training stability and final performance. (b) Impact of entropy and learning rate adaptation on exploration and convergence. (c) Performance under different mixing coefficients $p$ in flow policies.
@@
Entropy and Learning rate adaptation. Figure 4(b) illustrates the effect of entropy regularization and learning rate adaptation. Adaptive adjustment of the learning rate leads to faster and more stable convergence, while entropy regularization enhances exploration and yields higher returns. Empirical results confirm that both components, implemented as described in Section 4.3, substantially improve overall performance. In all tasks, we employ an adaptive learning rate scheme and include an entropy regularization term in the loss defined in Eq. (12).
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (62303319, 62406195), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), Shanghai Engineering Research Center of Intelligent Vision and Imaging. This work was also supported in part by computational resources provided by Fcloud CO., LTD.
@@ B.3 Hyperparameters
Learning rate for actor $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$
Learning rate for critic $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$
@@
Memory size N/A N/A $\times 10^{6} 1×1$$\times 10^{6}$ $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
@@
Memory size N/A N/A $\times 10^{6} 1×1$$\times 10^{6}$ $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
@@
Learning rate for actor $\times 10^{-4} 1×1$$\times 10^{-4}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$ $1\times 10^{-4}$
Learning rate for critic $\times 10^{-4} 1×1$$\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$ $1\times 10^{-4}$
@@
Memory size N/A N/A $\times 10^{6} 1×1$$\times 10^{6}$ $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
@@
Memory size N/A N/A $\times 10^{6} 1×1$$\times 10^{6}$ $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
@@
Learning rate for actor $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$
Learning rate for critic $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$
@@
Memory size N/A N/A $\times 10^{6} 1×1$$\times 10^{6}$ $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
@@
Learning rate for actor $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$
Learning rate for critic $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$
@@
Memory size N/A N/A $\times 10^{6} 1×1$$\times 10^{6}$ $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
@@
Learning rate for actor $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$ $5\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$
Learning rate for critic $\times 10^{-3} 1×1$$\times 10^{-3}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$
@@
Memory size N/A N/A $\times 10^{6} 1×1$$\times 10^{6}$ $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$
@@
Memory size N/A N/A $\times 10^{6} 1×1$$\times 10^{6}$ $\times 10^{6} 1×1$$\times 10^{6}$ $1\times 10^{6}$